(1 point) Two bonds, each with a face value of $18000, are redeemable at par in t-years and priced to yield y,-9%. Bond 1 has a coupon rate c4 = 11.4% and sells for $20225.65. Bond 2 has coupon rate c4-4.5% and sells for $ P. What is the value of P? Answer:$
Answer #1

Maturity=NPER(9%/4,11.4%/4*18000,-20225.65,18000)=28 or 7 years

P=PV(9%/4,7*4,4.5%*18000/4,18000)=13826.91